Use algebraic substitution to solve the following systems of equations
-2x+3y+5z=-21
-4z=20
6x-3y=0

Answer by Alan3354(69443)

equations
-2x+3y+5z=-21
-4z=20
6x-3y=0
----------
Step 1, Solve for z